# Energy extraction through Comisso-Asenjo process from a Kerr-Newman black hole in perfect fluid dark matter

### v1

@@ IV ENERGY EXTRACTION BY CAMISSO-ASENJO PROCESS
Assuming plasma fluid co-rotates around the equatorial plane of this rotating charged black hole in PFDM, the Keplarian velocity in Boyer-Lindquist coordinates can be obtained from
@@
Note that contravariant vectors in the ZAMO frames can be transformed to the Boyer-Lindquist coordinates by $\hat{b}^{0}=\alpha b^{0}$, and $\hat{b}^{i}=\sqrt{g_{ii}}b^{i}-\alpha\beta^{i}b^{0}$, while covectors transform by $\hat{b}_{0}=b_{0}/\alpha+\Sigma^{3}_{i=1}(\beta^{i}/\sqrt{g_{ii}}b_{i}$ and $\hat{b}_{i}=b_{i}/\sqrt{g_{ii}}$. With these transformations, the Keplarian velocity of the corotating bulk plasma in ZAMO frame can be found by $\hat{v}_{K}=\frac{d\hat{x}^{\phi}}{d\hat{t}}=\frac{\sqrt{g_{\phi\phi}}}{\alpha}{\Omega_{K}}-\beta^{\phi}$.
@@
where $\rho$, $\omega$, $U^{\mu}$, and $F^{\mu\nu}$ are the proper plasma pressure, enthalpy density, four-velocity, and the electromagnetic field tensor, respectively. To evaluate the energy measured by static observers at infinity, we construct a quantity, known as the “energy-at-infinity” density $e^{\infty}$, from $e^{\infty}\equiv n_{\mu}J^{\mu}=n_{\mu}T^{\mu\nu}\xi_{\nu}=-\alpha g_{\mu 0}T^{\mu 0}$, where $\xi_{\nu}=(\partial_{t},0,0,0)$ is the time-like Killing vector and $n_{\nu}$ is the unit vector normal to time-like hypersurfaces. Adopting Comisso and Asenjo’s assumption Comisso and Asenjo (2021), we consider the bulk plasma is incompressible and adiabatic, and the conversion of magnetic energy into plasma’s kinetic energy during magnetic reconnection is highly efficient. Under this condition, the energy density $e^{\infty}$ is mostly dominated by the hydrodynamic component while the electromagnetic component can be neglected, leading to an expression
@@
where $\hat{\gamma}$ is the outflow Lorentz factor and $\hat{v}^{\phi}$ is the azimuthal component of the outflow velocity measured by the ZAMO observer. In Comisso and Asenjo’s configuration, the current density is directed radially within the reconnection layer in the azimuthal direction. The outflow velocity and its associated Lorentz factor are expressed in terms of the plasma magnetization upstream of the reconnection layer $\sigma_{0}$
@@
where $\pm$ represents the corotating (+)(+)( + ) and counter-rotating ()(-)( - ) plasma flow relative to the black hole, and $\xi$ denotes the angle between the radial and azimuthal components of the outflow velocity. Combining equations 23 to 26, the final expression for the energy density at infinity per enthalpy $\epsilon^{\infty}=e^{\infty}/\omega$ leads to Comisso and Asenjo (2021)
@@
Similarly, as in the Penrose process, energy extraction for the rotating charged black hole in PFDM occurs when the decelerated plasma secures negative energy-at-infinity, while the accelerated plasma particles possess energy-at-infinity more than its rest mass and thermal energies. This requires the following conditions to be met Comisso and Asenjo (2021)
@@
where the polytropic index $\Gamma=4/3$ is assumed for relatively hot plasma.
IV.1 PARAMETER AND PHASE SPACE ANALYSIS
@@
Figure 3: Top figure: energy-at-infinity density per enthalpy, $\epsilon^{\infty}_{+}$ (blue lines) and $\epsilon^{\infty}_{-}$ (orange lines) as a function of plasma magnetization $\sigma_{0}$, for two $\lambda$ values with $(a,Q,\xi)$ fixed at $(1.0,1.0,0)$. Bottom figure: cutoff values of $\sigma_{0}$ as a function of $\lambda$ for three $Q$ values.
As we consider various combinations of the parameters ($a,\xi,\sigma_{0}$), it is important to note that energy extraction is favored by lower values of the orientation angle $\xi$ and higher values of $\sigma_{0}$, as implied by equation [reference] and demonstrated in the Kerr case [reference]. This condition is illustrated in the top figure of Fig. 3, where $\epsilon^{\infty}_{+}$ (for accelerating plasma) and $\epsilon^{\infty}_{-}$ (for decelerating plasma) are shown for two values of the dark matter parameter $\lambda$. In both cases, $\epsilon^{\infty}_{-}$ becomes more negative as $\sigma_{0}$ increases, but the cutoff value of $\sigma_{0}$, beyond which $\epsilon^{\infty}_{-}<0$ is secured, appears to decrease as $\lambda$ increases. The bottom figure of Fig. 3 shows that the dependence of the cutoff $\sigma_{0}$ on $\lambda$ is not always monotonic. Additionally, it indicates that $\lambda$ must remain below a certain critical value for a positive cutoff $\sigma_{0}$ to be possible. This is not surprising because, as previously discussed, such critical values for $\lambda$ can effectively influence horizons, which in turn determine the location of the reconnection X-point to occur within the ergoregion.
Figure 4: Energy-at-infinity density per enthalpy, $\epsilon^{\infty}_{+}$ (black lines) and $\epsilon^{\infty}_{-}$ (orange lines), as a function of the dark matter parameter $\lambda$ for varying $a$ with fixed $Q$ at 0.8 (top figure), and varying $Q$ with fixed $a$ at 0.8 (bottom figure). For both cases, we choose $\xi=0$ and $\sigma_{0}=100$.
@@
The impact of dark matter on ϵ±subscriptsuperscriptitalic-$\epsilon^{\infty}_{\pm}$ is further depicted in Fig 4 for different combinations of $a$ and $Q$. Both the top and bottom figures indicate that $\epsilon^{\infty}_{-}<0$ is met within a range of $\lambda$ values that widens as $Q$ decreases and $a$ increases. The values of $\lambda$ in these ranges increase monotonically as either $a$ or $Q$ increases.
Furthermore, regions of phase-space $(a,r/M)$ where energy extraction occurs $(\epsilon^{\infty}_{-}<0)$ are depicted in Fig 5 for different values of dark matter parameter. The increase in the area of this region as $\lambda$ decreases implies that lower values of $\lambda$ are favored in energy extraction for smaller values of $Q$, which is consistent with the results shown in Fig. 1 (e).
IV.2 ENERGY EXTRACTION RATE AND RECONNECTION EFFICIENCY
In this section, we continue to analyze the role of the parameter space on the energy extraction rate $P_{extr}$ and reconnection efficiency $\eta$. The energy extraction rate is determined by the amount of plasma with negative energy at infinity falling into the ergosphere and can be estimated as Comisso and Asenjo (2021)
@@ IV.2 ENERGY EXTRACTION RATE AND RECONNECTION EFFICIENCY
Figure 6: Extracted power per enthalpy $P_{extr}/\omega_{0}$ as a function of the X-point location with varying plasma magnetization (top figure) and varying orientation angle (bottom figure) with $(a,Q,\lambda)$ fixed at $(0.8,0.8,1.17)$.
We first show in Fig. 6 the extracted power per enthalpy, $P_{extr}/\omega_{0}$, as a function of the X-point location $r/M$ for different plasma magnetization values $\sigma_{0}\in\{10,10^{2},10^{3},10^{4},10^{5}\}$ with $(a,Q,\lambda,\xi)$ = $(0.8,0.8,1.17,0)$ in the top figure, and for different orientation angle values $\xi\in\{0,\pi/20,\pi/12,\pi/8,\pi/6\}$ with $(a,Q,\lambda,\sigma_{0})$ = $(0.8,0.8,1.17,100)$ in the bottom figure. Similar to the standard Kerr case Comisso and Asenjo (2021), the power extracted from the black hole favors high plasma magnetization and low orientation angle values until it drops off near the limiting circular orbits. For the Kerr-Newman black hole in the presence of PFDM, we also observe the increased sensitivity of the orientation angle on the extracted power and its drop-off limits. For the remainder of the analysis, we use the preferred values of $\xi=0$ and $\sigma_{0}=10^{4}$ based on these results.
@@
V Summary and Conclusions
As proposed by Comisso and Ansenjo Comisso and Asenjo (2021), magnetic reconnection can serve as an efficient mechanism for energy extraction from a rotating black hole in the ergosphere, where frame dragging causes anti-parallel magnetic field lines and a current sheet to form. Magnetic reconnection converts magnetic energy into plasma particle energy, producing rapid outflows; energy extraction occurs when the outflow moving opposite to the black hole’s rotation has negative energy-at-infinity, while the other outflow escapes to infinity with gained energy. This novel way of extracting a black hole’s rotational energy via rapid energy release during reconfiguration of the magnetic field has provided new insights into explaining bursty relativistic jets.
@@ V Summary and Conclusions
We first studied the geodesics of the black hole in question and looked for optimized cases where its ergoregion is reasonably sizable for energy extraction, and the innermost stable circular orbits (ISCO) reside within the outer boundary of the ergoregion. Since the metric components $g_{tt}$ and $g_{rr}$ involve a logarithmic term, analyzing the event horizons and circular geodesics becomes much more complex and can only be performed numerically. As summarized in Fig 1, we found that the event horizons depend on both the electric charge $Q$ and the dark matter parameter $\lambda$. Particularly, the outer horizon $r_{h+}$ initially decreases as $\lambda$ increases until reaching a critical value $\lambda_{c}$, beyond which $r_{h+}$ increases more rapidly as $\lambda$ grows. A similar non-monotonic trend holds for the variation of the size of ergoregion $r_{L+}-r_{h+}$ as a function of $\lambda$. We also found that the size of ergoregion can significantly increase at faster spins ($a>0.8$) as the dark matter parameter $\lambda$ decreases, given the electric charge stays within $Q\in[0.2,0.5]$. Despite the complex impact of parameters $(a,Q,\lambda)$ on the horizon structure, we showed that the condition required by magnetic reconnection $r_{isco}<r_{h+}$ is securely satisfied, as shown in Fig 2.
@@
Next, we evaluated the energy extraction conditions: energy-at-infinity per enthalpy for decelerating plasma $\epsilon^{\infty}_{-}<0$ and energy-at-infinity per enthalpy $\epsilon^{\infty}_{+}>0$ for various combinations of parameters involved $(a,Q,\lambda,\xi,\sigma_{0})$. Similar to the standard Kerr case, energy extraction in Kerr-Newman black hole in PFDM is also favored at higher values of plasma magnetization $\sigma_{0}$ and near-null orientation angle $\xi$. However, the presence of charge $Q$ and dark matter component $\lambda$ introduces additional features in determining the cutoff value of plasma magnetization that triggers energy extraction. As depicted in Fig 3, we found that this cutoff plasma magnetization can be lowered if the dark matter component increases or the charge decreases. Fig 4 shows that the permitted range of dark matter contribution for energy extraction to occur strongly depends on both the black hole spin and charge, benefited by lower values of charge. While keeping the spin and charge held at their favored values, a lower dark matter component results in greater regions in phase-space $(a,r/M)$ where energy extraction occurs.
@@
Lastly, we analyzed the power extracted per enthalpy $P_{ext}/\omega_{0}$ and reconnection efficiency $\eta$ as a function of X-point location for different values in the full parameter space $(a,Q,\lambda,\xi,\sigma_{0})$. Figs 6, 7(a) and 7(d) confirm that extracted power is still enhanced by high plasma magnetization, low orientation angle and high black hole spin, with the maximum value of $P_{ext}/\omega_{0}\approx 5$ and $\eta\approx 1.5$ achieved at near-unity spins comparable to the standard near-extremal Kerr case. Most interestingly, we found that the presence of charge and dark matter background could elevate $P_{ext}/\omega_{0}$ and $\eta$ to the same maximum values even when the black hole spin is not particularly high, as shown in panels (b)-(c) and (e)-(f) in Fig.7.
@@
An interesting avenue to pursue next is to explore how energy extraction via magnetic reconnection transpires in accretion disks surrounding black holes in conformal Weyl gravity. It is well known that black holes in this theory naturally encode dark matter and dark energy within their vacuum, without requiring adding dark matter and dark energy terms in the Lagrangian Mannheim and Kazanas (1989). Additionally, conformal Weyl gravity includes the Kerr spacetime as one of its vacuum solutions, allowing the CA framework to be naturally applied. Extending energy extraction studies to the more general conformal Weyl gravity, which enhances general relativity with an additional conformal symmetry, could reveal new and novel aspects of magnetic reconnection as an energy extraction mechanism from rotating black holes.
Acknowledgements.

### v2

@@ 4 ENERGY EXTRACTION BY CAMISSO-ASENJO PROCESS
Considering plasma fluid co-rotates around the equatorial plane of this rotating charged black hole in PFDM, the Keplarian velocity in Boyer-Lindquist coordinates can be obtained from
@@
Note that contravariant vectors in the ZAMO frames can be transformed to the Boyer-Lindquist coordinates by $\hat{b}^{0}=\alpha b^{0}$, and $\hat{b}^{i}=\sqrt{g_{ii}}b^{i}-\alpha\beta^{i}b^{0}$, while covectors transform by $\hat{b}_{0}=b_{0}/\alpha+\Sigma^{3}_{i=1}(\beta^{i}/\sqrt{g_{ii}})b_{i}$ and $\hat{b}_{i}=b_{i}/\sqrt{g_{ii}}$. With these transformations, the Keplarian velocity of the corotating bulk plasma in ZAMO frame can be found by $\hat{v}_{K}=\frac{d\hat{x}^{\phi}}{d\hat{t}}=\frac{\sqrt{g_{\phi\phi}}}{\alpha}{\Omega_{K}}-\beta^{\phi}$.
@@
where $\rho$, $\omega$, $U^{\mu}$, and $F^{\mu\nu}$ are the proper plasma pressure, enthalpy density, four-velocity, and the electromagnetic field tensor, respectively. Note that we use $T^{\mu\nu}_{b}$ here to distinguish it from the energy-momentum tensor $T^{s}_{\mu\nu}$ that sources the black hole. We assume the spacetime is fixed and the bulk plasma does not affect the black hole’s charge $Q$. Therefore, the electromagnetic field tensor $F_{\mu\nu}$ in the above equation refers to the contribution from the bulk plasma only.
Similar to the Penrose process, energy extraction via magnetic reconnection to occur at the ergosphere, the accelerated plasma escaping the black hole attains positive energy, while the decelerated plasma absorbed by the black hole must hold negative energy. To evaluate the energy measured by static observers at infinity, we construct a quantity, known as the “energy-at-infinity” density $e^{\infty}$
$e^{\infty}\equiv n_{\mu}J^{\mu}=n_{\mu}T^{\mu\nu}_{b}\xi_{\nu}=-\alpha g_{\mu 0}T^{\mu 0}_{b}=\alpha\hat{e}+\alpha\beta^{\phi}\hat{P}^{\phi},$ (25)
where $\xi_{\nu}=(\partial_{t},0,0,0)$ is the time-like Killing vector and $n_{\nu}$ is the unit vector normal to time-like hypersurfaces. The total energy density $\hat{e}$ and azimuthal component of the momentum density $\hat{P}^{\phi}$ are given as
$\begin{split}\hat{e}&=\omega\hat{\gamma}^{2}-p+\frac{\hat{B}^{2}+\hat{E}^{2}}{2},\\ \hat{P}^{\phi}&=\omega\hat{\gamma}^{2}\hat{v}^{\phi}+(\hat{B}\times\hat{E})^{\phi},\end{split}$ (26)
where the outflow Lorentz factor $\hat{\gamma}=\hat{U}^{0}=\sqrt{1-\Sigma^{3}_{i=1}(d\hat{v}^{i})^{2}}$ and $\hat{v}^{\phi}$ is the azimuthal component of the outflow velocity measured by the ZAMO observer. The electric and magnetic field components are defined by
$\begin{split}\hat{B}^{i}&=\frac{1}{2}\epsilon^{ijk}F_{jk},\\ \hat{E}^{i}&=\eta^{ij}F_{j0}.\end{split}$ (27)
The energy density at infinity can be further rewritten in terms of its hydrodynamic and electromagnetic components: $e^{\infty}=e^{\infty}_{hyd}+e^{\infty}_{em}$, expressed as
$\begin{split}e^{\infty}_{hyd}&=\alpha\hat{e}_{hyd}+\alpha\beta^{\phi}\omega\hat{\gamma}^{2}\hat{v}^{\phi},\\ e^{\infty}_{em}&=\alpha\hat{e}_{em}+\alpha\beta^{\phi}(\hat{B}\times\hat{E})_{\phi},\end{split}$ (28)
where $\hat{e}_{hyd}=\omega\hat{\gamma}^{2}-p$ and $\hat{e}_{em}=(\hat{B}^{2}+\hat{E}^{2})/2$ are the energy densities of the hydrodynamic and electromagnetic fields in the ZAMO frame. Adopting Comisso and Asenjo’s assumption [46], we consider the bulk plasma is incompressible and adiabatic, and the conversion of magnetic energy into plasma’s kinetic energy during magnetic reconnection is highly efficient. Under this condition, the energy density at infinity $e^{\infty}$ is predominantly dominated by the hydrodynamic component, with the electromagnetic contribution being negligible, leading to an expression
@@
Figure 3: Left figure: energy-at-infinity density per enthalpy, $\epsilon^{\infty}_{+}$ (blue lines) and $\epsilon^{\infty}_{-}$ (orange lines) as a function of plasma magnetization $\sigma_{0}$, for two $\lambda$ values with $(a,Q,\xi)$ fixed at $(1.0,1.0,0)$. Right figure: cutoff values of $\sigma_{0}$ as a function of $\lambda$ for three $Q$ values.
In Comisso and Asenjo’s configuration, the current density is directed radially within the reconnection layer in the azimuthal direction. The outflow velocity and its associated Lorentz factor are expressed in terms of the plasma magnetization upstream of the reconnection layer $\sigma_{0}$
@@
where $\pm$ represents the corotating (+)(+)( + ) and counter-rotating ()(-)( - ) plasma flow relative to the black hole, and $\xi$ denotes the angle between the radial and azimuthal components of the outflow velocity. Combining equations 29 to 32, the final expression for the energy density at infinity per enthalpy $\epsilon^{\infty}=e^{\infty}/\omega$ leads to [46]
@@
Similarly, as in the Penrose process, energy extraction for the rotating charged black hole in PFDM occurs when the decelerated plasma secures negative energy-at-infinity, while the accelerated plasma particles possess energy-at-infinity more than its rest mass and thermal energies. This requires the following conditions to be met [46]
@@
where the polytropic index $\Gamma=4/3$ is assumed for relativistic hot plasma.
Figure 4: Energy-at-infinity density per enthalpy, $\epsilon^{\infty}_{+}$ (black lines) and $\epsilon^{\infty}_{-}$ (orange lines), as a function of the dark matter parameter $\lambda$ for varying $a$ with fixed $Q$ at 0.8 (left figure), and varying $Q$ with fixed $a$ at 0.8 (right figure). For both cases, we choose $\xi=0$ and $\sigma_{0}=100$.
4.1 PARAMETER AND PHASE SPACE ANALYSIS
@@
As we consider various combinations of the parameters ($a,\xi,\sigma_{0}$), it is important to note that energy extraction is favored by lower values of the orientation angle $\xi$ and higher values of $\sigma_{0}$, as implied by equation 33 and demonstrated in the standard Kerr case [46]. This condition is illustrated in the left figure of Fig. 3, where $\epsilon^{\infty}_{+}$ (for accelerating plasma) and $\epsilon^{\infty}_{-}$ (for decelerating plasma) are shown for two values of the dark matter parameter $\lambda$. In both cases, $\epsilon^{\infty}_{-}$ becomes more negative as $\sigma_{0}$ increases, but the cutoff value of $\sigma_{0}$, beyond which $\epsilon^{\infty}_{-}<0$ is secured, appears to decrease as $\lambda$ increases. The right figure of Fig. 3 shows that the dependence of the cutoff $\sigma_{0}$ on $\lambda$ is not always monotonic. Additionally, it indicates that $\lambda$ must remain below a certain critical value for a positive cutoff $\sigma_{0}$ to be possible. This is not surprising because, as previously discussed, such critical values for $\lambda$ can effectively influence horizons, which in turn determine the location of the reconnection X-point to occur within the ergoregion.
The impact of dark matter on ϵ±subscriptsuperscriptitalic-$\epsilon^{\infty}_{\pm}$ is further depicted in Fig 4 for different combinations of $a$ and $Q$. Both the left and right figures indicate that $\epsilon^{\infty}_{-}<0$ is met within a range of $\lambda$ values that widens as $Q$ decreases and $a$ increases. The values of $\lambda$ in these ranges increase monotonically as either $a$ or $Q$ increases. Fig 5 illustrates the strong influence of the dark matter parameter on the regions of phase-space $(a,r/M)$ where energy extraction occurs $(\epsilon^{\infty}_{-}<0)$. It shows a clear trend that the phase-space region shifts to a larger radial distance as $\lambda$ increases and this region also expands as $\lambda$ decreases. The key point to take away is that the phase-space region can extend to a larger radial location where energy extraction favors due to the combined effect of the black hole charge and dark matter parameter.
@@
4.2 ENERGY EXTRACTION RATE AND RECONNECTION EFFICIENCY
In this section, we continue to analyze the role of the parameter space on the energy extraction rate $P_{extr}$ and reconnection efficiency $\eta$. The energy extraction rate is determined by the amount of plasma with negative energy at infinity falling into the ergosphere and can be estimated as [46]
@@ 4.2 ENERGY EXTRACTION RATE AND RECONNECTION EFFICIENCY
Figure 6: Extracted power per enthalpy $P_{extr}/\omega_{0}$ as a function of the X-point location with varying plasma magnetization (left figure) and varying orientation angle (right figure) with $(a,Q,\lambda)$ fixed at $(0.8,0.8,1.17)$.
We first show in Fig. 6 the extracted power per enthalpy, $P_{extr}/\omega_{0}$, as a function of the X-point location $r/M$ for different plasma magnetization values $\sigma_{0}\in\{10,10^{2},10^{3},10^{4},10^{5}\}$ with $(a,Q,\lambda,\xi)$ = $(0.8,0.8,1.17,0)$ in the left figure, and for different orientation angle values $\xi\in\{0,\pi/20,\pi/12,\pi/8,\pi/6\}$ with $(a,Q,\lambda,\sigma_{0})$ = $(0.8,0.8,1.17,100)$ in the right figure. Similar to the standard Kerr case [46], the power extracted from the black hole favors high plasma magnetization and low orientation angle values until it drops off near the limiting circular orbits. For the Kerr-Newman black hole in the presence of PFDM, we also observe the increased sensitivity of the orientation angle on the extracted power and its drop-off limits. For the remainder of the analysis, we use the preferred values of $\xi=0$ and $\sigma_{0}=10^{4}$ based on these results.
@@
5 Summary and Conclusions
As proposed by Comisso and Ansenjo [46], magnetic reconnection can serve as an efficient mechanism for energy extraction from a rotating black hole in the ergosphere, where frame dragging causes anti-parallel magnetic field lines and a current sheet to form. Magnetic reconnection converts magnetic energy into plasma particle energy, producing rapid outflows; energy extraction occurs when the outflow moving opposite to the black hole’s rotation has negative energy-at-infinity, while the other outflow escapes to infinity with gained energy. This novel way of extracting a black hole’s rotational energy via rapid energy release during reconfiguration of the magnetic field has provided new insights into explaining bursty relativistic jets.
@@ 5 Summary and Conclusions
We first studied the geodesics of the black hole in question and looked for optimized cases where its ergoregion is reasonably sizable for energy extraction. Since the metric components $g_{tt}$ and $g_{rr}$ involve a logarithmic term, analyzing the event horizons and circular geodesics becomes much more complex and can only be performed numerically. As summarized in Fig 1 and Fig 2, we found that the event horizons and circular photo orbits strongly depend on both the electric charge $Q$ and the dark matter parameter $\lambda$. Particularly, the outer horizon $r_{h+}$ and co-rotating photon orbit $r_{p}$ initially decreases as $\lambda$ increases until reaching a critical value $\lambda_{c}$, beyond which $r_{h+}$ and $r_{p}$ increase more rapidly as $\lambda$ grows. A similar non-monotonic trend holds for the variation of the size of ergoregion $r_{L+}-r_{h+}$ as a function of $\lambda$. We also found that the size of ergoregion can significantly increase at faster spins ($a>0.8$) as the dark matter parameter $\lambda$ decreases, given the electric charge stays within $Q\in[0.2,0.5]$.
@@
Next, we evaluated the energy extraction conditions: energy-at-infinity per enthalpy for decelerating plasma $\epsilon^{\infty}_{-}<0$ and energy-at-infinity per enthalpy $\epsilon^{\infty}_{+}>0$ for various combinations of parameters involved $(a,Q,\lambda,\xi,\sigma_{0})$. Similar to the standard Kerr case, energy extraction in Kerr-Newman black hole in PFDM is also favored at higher values of plasma magnetization $\sigma_{0}$ and near-null orientation angle $\xi$. However, the presence of charge $Q$ and dark matter component $\lambda$ introduces additional features in determining the cutoff value of plasma magnetization that triggers energy extraction. As depicted in Fig 3, we found that this cutoff plasma magnetization can be lowered if the dark matter component increases or the charge decreases. Fig 4 shows that the permitted range of dark matter contribution for energy extraction to occur strongly depends on both the black hole spin and charge. Additionally, a lower dark matter component results in greater regions in phase-space $(a,r/M)$, while increasing the dark matter parameter can also extend the phase-space region to a larger radial distance favored by energy extraction.
@@
Lastly, we analyzed the power extracted per enthalpy $P_{ext}/\omega_{0}$ and reconnection efficiency $\eta$ as a function of X-point location for different values in the full parameter space $(a,Q,\lambda,\xi,\sigma_{0})$. Figs 6, 7(a) and 7(d) confirm that extracted power is still enhanced by high plasma magnetization, low orientation angle and high black hole spin, with the maximum value of $P_{ext}/\omega_{0}\approx 5$ and $\eta\approx 1.5$ achieved at near-unity spins comparable to the standard near-extremal Kerr case. Most interestingly, we found that the presence of charge and dark matter background could elevate $P_{ext}/\omega_{0}$ and $\eta$ to the same maximum values even when the black hole spin is not particularly high, as shown in panels (b)-(c) and (e)-(f) in Fig.7. It relaxes the stringent requirements found in other types of rotating black holes [47, 48, 49, 50, 51, 52, 53, 54, 55, 56], where near-extremal spin is typically necessary to achieve a comparable level of extracted power and reconnection efficiency.
@@
An interesting avenue to pursue next is to explore how energy extraction via magnetic reconnection transpires in accretion disks surrounding black holes in conformal Weyl gravity. It is well known that black holes in this theory naturally encode dark matter and dark energy within their vacuum, without requiring adding dark matter and dark energy terms in the Lagrangian [83]. Additionally, conformal Weyl gravity includes the Kerr spacetime as one of its vacuum solutions, allowing the CA framework to be naturally applied. Extending energy extraction studies to the more general conformal Weyl gravity, which enhances general relativity with an additional conformal symmetry, could reveal new and novel aspects of magnetic reconnection as an energy extraction mechanism from rotating black holes.
Acknowledgments